OpenStudy (anonymous):

Need help with solving this and show steps please sqrt{10-3x=x

OpenStudy (anonymous):

\[\sqrt{?10-3x=x}\]

OpenStudy (turingtest):

\[\sqrt{10-3x}=x\] you mean, right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

square both sides: \[10-3x=x^2\] rewrite as \[x^2+3x-10=0\] factor if possible (here it is, otherwise you need the quadratic equation) \[(x+5)(x-2)=0\] x={-5,2} get it?

OpenStudy (anonymous):

yes! thank you

OpenStudy (turingtest):

no prob. that'll be one medal please :)

OpenStudy (anonymous):

I didnt know you can give medals

OpenStudy (anonymous):

what is the meaning of that?

OpenStudy (anonymous):

\[7x-4\le 4+5x\]